OpenStudy (anonymous):

Find the quotient –7 ÷ 14/3

OpenStudy (anonymous):

a) -2/3 b) -98/3 c) -1/2 d) -3/2

OpenStudy (anonymous):

sometimes its easier to multiply by the reciprocal when dividing: \[-7 \div \frac{ 14 }{ 3 } = -7 * \frac{ 3 }{ 14 } = \frac{ -7 }{ 1 } * \frac{ 3 }{ 14 }\]